OpenStudy (anonymous):

Does anyone remember how to find slope? Im like totally brain washed and can someone help me find slope of the coordinates (-2,-3), (1,1)

OpenStudy (helder_edwin):

\[\large m=\frac{y_1-y_2}{x_1-x_2} \]

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

aha thank you @helder_edwin @jigglypuff314

OpenStudy (amigatour):

The formula is m=(y2-y1)/(x2-x1) y2=1 y1=-3 x2= 1 x1= -2 m=(1-(-3))/(1-(-2)) m=(1+3)/(1+2) m=4/3

OpenStudy (anonymous):

ok thx

OpenStudy (amigatour):

yw

OpenStudy (amigatour):

The other way of finding it is by graphing the given ordered pairs.

OpenStudy (anonymous):

okie dokie :)

OpenStudy (amigatour):